OpenStudy (erich):

2x^2-x-3=3 Solve for X

OpenStudy (erich):

i did (2x-3)(x+1)=0

OpenStudy (saifoo.khan):

I gave you medal bcz u r right. :D

OpenStudy (erich):

thanks, but how do I solve from there? is one answer 1.5 and the other -1?

OpenStudy (saifoo.khan):

RIght.

OpenStudy (saifoo.khan):

2x - 3 = 0 & (x + 1) = 0

OpenStudy (anonymous):

hey is this the question \[2x^2-x-3=3\]

OpenStudy (erich):

no

OpenStudy (erich):

it equals 0

OpenStudy (anonymous):

ok \[2x^2-x-3=0\] \[2x^2-3x+2x-3=0\] \[x(2x-3)+1(2x-3)=0\] \[(x+1)(2x-3)=0\] either x+1 = 0 or 2x-3 = 0 x=-1 or x=-3/2